### v1

# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

Abstract.
Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, holds promise for empowering individuals and transforming healthcare delivery. However, barriers to its adoption remain underexplored in the Indian context. This study investigates PHI adoption among Indian users and stakeholders using a multi-method approach. An awareness survey (n = 87) examined the usage of wearables and general PHI engagement, followed by semi-structured interviews (n = 22) that explored motivations, usage patterns, and health information sources. Qualitative analysis revealed that while PHI is valued for health monitoring and shared/collective care, its adoption is hindered by factors such as low health literacy, usability challenges, and mistrust in digital health platforms. Further stakeholder interviews and co-design workshops informed the development of a Figma-based prototype, which was evaluated for usability. Based on these findings, we offer design recommendations for an integrated, user-controlled PHI platform featuring accessible analytics and verifiable health information. Our insights highlight the socio-technical challenges of PHI adoption in India and underscore the need for reliable, user-centric solutions to support proactive healthcare.
1. Introduction
Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling active self-management. Globally, PHI leverages technologies—from smartphone applications to wearable devices such as smartwatches and fitness trackers—facilitating continuous health monitoring (Shajari et al., 2023). Factors like high health literacy, technology familiarity, and a proactive healthcare ethos drive PHI adoption in the Western context (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024). Emerging evidence (Chmielewski and Meyer, 2024) in the Western context suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care—seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
In India, however, infrastructural, sociocultural, and technological challenges intersect and constrain PHI’s potential. Digital health tools, including wearables, are often perceived as fashion accessories or status symbols rather than integrated components of comprehensive health management (Misra et al., 2023). Moreover, cultural practices, family dynamics, and skepticism towards advanced technologies, such as AI-driven solutions, further influence PHI usage and adoption (Hoque and Bao, 2015; Zhang et al., 2022; Creaser et al., 2022; Patil et al., 2022; Reis et al., 2024; Kapania et al., 2022; Wisetsri and Vijai, 2021). Given the cultural dynamics of the Indian context, a significant gap in its adoption and use can be filled by examining and understanding how PHI is currently integrated and utilized in daily health practices in India and shared with family and loved ones.
To address these challenges, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach. We conducted a survey (n=87) to assess awareness and usage of smart wearables, followed by interviews (n=22) exploring user motivations, usage patterns, and health information sources. Additional interviews with key stakeholder pairs—couples, intergenerational family groups, and healthcare providers—and co-design workshops (n=12) employing card sorting examined shared care practices and data-sharing preferences. Further, we evaluated a figma prototype to understand its usability and get design recommendations from the participants on what, how, and when they would like to share their health information/analytics with different stakeholders.
Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India. Specifically, we ask: (1) How do Indians adopt, integrate, and perceive PHI in their health management? (2) What dictates PHI’s effective utilization and perceived utility, and how can we improve them?
Our findings offer a deeper understanding of PHI challenges and opportunities in India and provide a Proof-of-Concept evaluation to strengthen our recommendations further. Through this work, we lay the groundwork for a more reliable and integrated PHI system supporting personal and shared care management, culturally appropriate to the Indian context and the global south.
2. Related Work
The body of work on personal health informatics within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019). These factors have propelled research toward innovative health management solutions and broadened the adoption of PHI systems. The rapid increase in chronic health conditions and an expanding digital landscape make personal health informatics especially critical in India. This work aims to contextualize the evolution of PHI within the Indian healthcare landscape, addressing its unique challenges and opportunities. This section examines the existing scholarship through two critical lenses: (1) personal health informatics in HCI and digital health tracking and (2) the barriers and enablers of PHI in the Indian context. These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural health practices.
2.1. Personal Health Informatics in HCI
Personal health informatics (PHI) within HCI has evolved from individual tracking (Rooksby et al., 2014) to encompassing broader ecological models (Epstein et al., 2015) that account for social, structural, and cultural factors. Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts. These frameworks have been extensively applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022), emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing.
Wearable technologies and smartphone applications are at the forefront of PHI in various contexts, like sports, managing chronic conditions, and broader long-term health monitoring initiatives (Tholander and Nylander, 2015; Klasnja et al., 2011; Poole, 2013; Nunes et al., 2015; Zegura et al., 2018; Blandford, 2019; Gooch et al., 2021; Fritz et al., 2014). Prior research on wearable use for PHI has also examined why people use and abandon smart devices (Lazar et al., 2015) and how it differs among people with prior self-tracking experience (Rapp and Cena, 2016). Studies on PHI often highlight the continuous capture and analysis of biosignals for precise intervention, particularly in chronic disease management, such as multiple sclerosis, chronic fatigue, and hypertension (Stuart et al., 2022; Jones et al., 2021; Tong et al., 2019).
While proactive use of PHI systems exists for critical use cases, such as emergencies (Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024), fall detections(Delahoz and Labrador, 2014; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022), and medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), we are yet to see them used for a broader PHI in the Indian context.
PHI in HCI is not limited to wearables and trackers; it also uses other modalities to integrate into an individual’s wellbeing. These diverse modalities represent different engagement strategies, from gamification in AR for improving physical activity (Zhao et al., 2024) using virtual pets to fostering social interaction through community displays (Gooch et al., 2021). Beyond physical and social wellbeing, wearables increasingly support mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023). Despite this, the literature (Lee et al., 2022) often emphasizes specialized healthcare needs, such as technology readiness and e-health literacy, while general proactive health tracking remains underexplored in HCI, particularly in India. Our work adopts an ecological perspective to examine intrinsic and extrinsic factors influencing PHI adoption, bridging the gap between research on wearables and broader proactive health management strategies.
2.2. Barriers and Enablers of PHI in the Indian Context
Kabir and Wiese’s meta-synthesis provides a broader understanding of the general global barriers and facilitators for personal informatics systems (Kabir and Wiese, 2023). Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges. Limited digital and health literacy among marginalized populations poses significant challenges. Older adults often face steep learning curves when adopting digital health technologies (Cherian et al., 2024), while children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023). Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement (Thakkar et al., 2022; Okolo et al., 2024; Lee et al., 2024b; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap.
Efforts to improve engagement through public health information datafication have shown promise in enhancing accessibility and comprehension (Huang et al., 2022; Pai and Alathur, 2020). However, implementing these strategies in India demands policy-level changes to ensure equitable access and integration with local health systems. National healthcare initiatives—such as Ayushman Bharat Health Account (ABHA) and expanding telemedicine services—have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023). Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024). However, sustaining long-term engagement remains challenging, as users frequently cite usability issues and information overload as key reasons for abandonment (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024).
Apart from the Individual and Public/Infrastructural challenges, cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023). Collectivist norms and uncertainty avoidance play a significant role in shaping attitudes toward digital health technologies (Karnatak et al., 2023; Zhang et al., 2022). The perceived economic value of health technologies is often underappreciated unless they provide critical utility to users through financial implications, for example, in managing chronic conditions (Singh et al., 2022b). While trust and perception issues persist, privacy concerns, data accuracy doubts, and discomfort with wearables further inhibit adoption (Shajari et al., 2023; Seelam et al., 2024).
While these multi-faceted barriers exist in the Indian context, our study aims to reconfirm these, identify more, and develop strategies to improve health literacy, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization. Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies.
3. Methodology
To examine the factors influencing the adoption of PHI in India, we conducted a study starting in January 2024, following approval from the Institute’s IRB. Our investigation began with a thorough review of the existing literature to understand the PHI usage and adoption landscape. It focused on the barriers and drivers for its adoption in personal health management and how its use (or the lack of it) is integrated into health management practices. We then designed a mixed-methods approach comprising surveys, interviews, and co-design workshops to capture the diverse perspectives of users. After analyzing all participant data and their inputs, we evaluated a rapid figma prototype of sharing preferences to further reinforce a unique finding.
3.1. Participants
The study recruited participants from across India, aiming to capture a diverse demographic range in age, gender, and professional background. To ensure a representative sample (especially those who can purchase and use even affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale by Saleem et al. (Saleem and Jan, 2021). We employed a combination of purposive and snowball sampling methods to identify and engage participants. The sample included a variety of professions, such as IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society. This approach enabled us to gather insights from participants with differing levels of familiarity and engagement with wearable health technologies. Prior to participation, all individuals were provided with a clear explanation of the study’s purpose, and informed consent was obtained. Demographic details for the surveys and interview participants are summarized in Table 1, offering a comprehensive overview of the sample’s composition.
Study Component | Total | Gender | Age | Education | |||||||
Male | Female | Prefer Not | Range | Mean | High | Bachelors | Advanced | ||||
to Say | School | Degrees | |||||||||
PHI awareness survey | 87 | 50 | 36 | 1 | 18-83 | 32 | 6 | 31 | 50 | ||
34 | 20 | 14 | 0 | 18-62 | 32 | 1 | 16 | 17 | ||
interviews | 22 | 16 | 8 | 0 | 18-62 | 28 | 5 | 6 | 11 | ||
co-design workshops | 12 | 6 | 6 | 0 | 27-40 | 33 | 0 | 8 | 4 | ||
114 | 44 | 70 | 0 | 19-75 | 34 | 3 | 57 | 54 |
[The table listing participant statistics from various study components]Table 1 lists the ranges and mean of age, gender, and educational background for the three components of the study.
3.2. Study Design
3.2.1. Survey Design
The initial PHI awareness survey was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants. The survey included closed-ended and open-ended questions designed to capture a broad spectrum of wearable health technology use data. The survey comprised a total of 29 questions, including 8 questions on demographic information and 8 open-ended questions (6 for users and 2 for non-users). The closed-ended questions focused on collecting demographic information (age, gender, occupation), the type of wearable devices used for PHI, frequency of use, and specific health metrics tracked (such as steps, heart rate, and sleep patterns). The open-ended questions focused on aspects of trust and usage, such as (1)Please explain briefly if you have ever changed the health wearable you use and why. (2) What do you understand about the health data/analysis from your wearable? (3) Do you trust the accuracy of the wearable health data? (4) Has the history from your health wearable ever helped you with doctor visits? The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives. These responses were crucial for identifying recurring themes and patterns in the qualitative analysis. The survey remained open until saturation was observed in the responses.
The PHI awareness survey had a total of 99 responses, of which 87 were complete and were used for further analysis. The demographics had a mean age of 32 and a maximum age of 83, which included 50 males, 36 females, and 1 participant who preferred not to disclose their gender. Among the complete responders, there were 55 users and 32 non-users. The majority (85%) of the survey participants were between 18 and 44 years old. The diversity of participants in age, gender, profession, and education provided a robust dataset for analyzing broader adoption/use/behavior patterns and informed the qualitative exploration in the subsequent interviews.
The concluding survey on health-sharing features was designed to assess both awareness and utilization of these features among Android and iOS users. It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each subgroup (Apple users, Apple non-users, Android users, Android non-users). Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions. The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and—when applicable—the frequency of their usage. Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses. The survey was closed upon reaching saturation of responses.
The health sharing feature survey was completed by 114 participants, with a mean age of 34 and a maximum age of 75, which included 44 males and 70 females. Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed. This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms, especially with the Android users (who are % of users in India).
ID | Age | Gender | Education | Wearable Usage | Was Gifted | Abandonment | Health Analytics |
P1 | 19 | Male | High School | Health | Yes | Still Using | Advanced |
P2 | 25 | Female | Advanced Degree | Non-User | N/A | N/A | N/A |
P3 | 33 | Female | Advanced Degree | Health | Yes | Still Using | Advanced |
P4 | 33 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P5 | 29 | Female | Bachelors | Utility | No | Still Using | N/A |
P6 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P7 | 32 | Male | Doctorate | Health/Utility | No | Still Using | Basic |
P8 | 26 | Male | Advanced Degree | Health | No | Still Using | Basic |
P9 | 29 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P10 | 62 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P11 | 34 | Male | Bachelors | Health/Utility | No | Stopped Using | Advanced |
P12 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P13 | 29 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Advanced |
P14 | 20 | Male | High School | Non-User | N/A | N/A | N/A |
P15 | 25 | Female | Advanced Degree | Fashion/Health | Yes | Still Using* | Basic |
P16 | 26 | Female | Advanced Degree | Health/Fashion | No | Still Using | Basic |
P17 | 26 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Basic |
P18 | 18 | Male | High School | Health/Utility | No | Still Using | Advanced |
P19 | 36 | Male | Advanced Degree | Health/Utility | No | Stopped Using** | Advanced |
P20 | 29 | Male | Advanced Degree | Non-User | N/A | N/A | N/A |
P21 | 27 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P22 | 27 | Female | Advanced Degree | Health/Utility | Yes | Stopped Using | Basic |
[The table listing participant demographics]Table 2 lists details of age, gender, education, usage/non-usage, level of health analytics, and Usage Abandonment for each participant by the uniquely assigned participant ID from the semi-structured Interviews.
3.2.2. Interview Design
The secondary component consisted of semi-structured interviews with 22 participants, all of whom were selected from the survey pool. The interviews were conducted based on the availability of the participants and were conducted until response saturation was reached on the emerging themes. Participants were chosen to reflect on various experiences with wearable technology for their PHI. The interview group included 16 males and 8 females, with a mean age of 28, from diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees. The detailed demographics are presented in Table 2. The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or stopped using their wearable devices. This diversity in the interview sample allowed for a comprehensive exploration of the different attitudes and behaviors toward wearable health technologies for PHI.
The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables. The interviews also explored participants’ broader health management practices and how wearables for PHI fit within this ”ecology of tracking.” This approach allowed the researchers to capture rich, contextual data that provided insights beyond what could be obtained from the survey alone.
ID | Age | Gender | Education | No. of Children |
CP1 | 34 | Female | Advanced Degree | 0 |
CP2 | 35 | Male | Bachelors | 0 |
CP3 | 33 | Female | Advanced Degree | 0 |
CP4 | 32 | Male | Bachelors | 0 |
CP5 | 27 | Female | Bachelors | 0 |
CP6 | 30 | Male | Bachelors | 0 |
CP7 | 33 | Male | Advanced Degree | 1 |
CP8 | 31 | Female | Advanced Degree | 1 |
CP9 | 29 | Female | Bachelors | 1 |
CP10 | 31 | Male | Bachelors | 1 |
CP11 | 38 | Female | Bachelors | 2 |
CP12 | 40 | Male | Bachelors | 2 |
[The table listing participant demographics of card sorting activity.]Table 3 lists details of age, gender, education, and number of children for each participant by the uniquely assigned participant ID from the co-design workshops.Advanced degree includes graduate degrees (like M.A., L.L.M., M.Sc., MS, MTech., M.Eng., MBA, CA, MD, MPhil, PhD, DPhil, etc.)
[Sample Card Sorting Activity Screens.]The screenshot shows the interface of the card sorting activity used by couples to group cards comprising of different health data (such as heart rate, SpO2, menstrual cycles, etc.) and aspects of sharing (such as control, privacy, and storage, etc.) to be sorted into categories of various stakeholder groups (such as parents, siblings, friends, doctors, etc.)
3.2.3. Co-design Workshops
One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information. This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI, such as Spouses, Parents/Adult Children, and Doctors, through the co-design workshops. To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6), Older Parent/Adult Children pairs (4), and doctors (4 from 3 specialties). To further identify the sharing preferences, we conducted a card sorting activity using kartSort (just using the couples as older participants struggled with the activity in the digital format and in-person activity was unfeasible as they lived across India) where they moved around different data/information cards to different stakeholders as categories. A sample screen from the card sorting activity is presented in Figure 1. We recruited six couples to reach saturation in understanding the variation of preferences among couples with and without a child (or children). The detailed demographics of the card sorting pairs are presented in Table 3.
[Sample Rapid Prototyping Screens.]The sample screenshot of the Figma prototype shows different screen flows in the prototype for selecting sharing preferences and approving requests for sharing among family members.
3.2.4. Rapid Prototyping and Evaluation
After the co-design workshops, we developed a medium fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members. This process helped us understand the ideal sharing preferences of individuals with various stakeholders involved. Figure 2 gives a sample screenshot of the prototype. The figma prototype was evaluated informally by sharing it with the participants and eliciting a qualitative evaluation. The unanimous positive takeaway from the participants’ feedback included its simplicity, modularity, and ease of control. Apart from these, there were individual features that the participant appreciated and provided further recommendations, which are discussed in the findings of this work. It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness, usage/non-usage of the same. The details of those observations are also presented in the findings, and the demographics were presented in 1.
Themes | Sub-Themes | Main Codes |
Collective health management | Use as a care device | Tracking heart rate & step count |
through sharing practices | Using because it was a gift | |
Track parents/partners health metrics | ||
Use as a health indicator | Use to monitor stress | |
Use to check critical conditions | ||
Step tracking as primary use | ||
Sleep tracking awareness | ||
Specific health tracking need | ||
(e.g., menstrual cycle) | ||
Digital PHI as a motivational tool | Perception on Wearable | Handy tool for tracking and utility |
for proactive health management | Useful motivational tool | |
Perception affecting degree of use | ||
Perceived low utility | ||
Trust in Data & Technology | Conditional trust in certain metrics | |
Acceptance of minor inaccuracies | ||
Useful as an approximate indicator | ||
Validation through comparison | ||
Low medical literacy on wearable data | ||
Navigating health information | Technical and Logistical Issues | Analytics not easy to understand |
in a fragmented digital ecosystem | Device management issues | |
Device incompatibilities | ||
Device change issues | ||
Buying a wearable is an investment | ||
Lifestyle and Physical Barriers | Lack of preferred modality of wearable | |
Rashes from the strap | ||
Minimal impact on doctor visits | ||
Physical discomfort on consistent use | ||
Preference for traditional methods | Trust on Body feeling | Will go to the doctor |
and perception of non-necessity | if I feel something wrong | |
Trust in health lifestyle/habits | ||
Perceptive lack of current need | ||
Reluctance to trust something new | ||
Trust in Doctors and | Will choose doctor for critical issues | |
digital PHI overwhelming | Do not know enough, doctors reliable | |
Wearable use overwhelming | ||
reactive care is enough | ||
Difference between body feeling | ||
and wearable metrics |
[Thematic Analysis table.]Table 4 lists the main themes, sub-themes, and the main codes obtained from the reflexive thematic analysis from the interview transcripts and the field notes taken during the interviews.
3.3. Data Analysis
We employed a combination of mixed methods to analyze the data. The survey responses were processed using descriptive statistics, where we calculated frequencies and percentages to assess patterns in the use and adoption of wearable health technologies for PHI, along with participants’ trust levels and perceived usefulness of these devices. To explore associations between demographic characteristics (such as age and gender) and usage patterns, we conducted cross-tabulations to identify significant relationships. We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, and field notes for the qualitative analysis. The analysis proceeded in several stages. Initially, the interview transcripts and field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns. Subsequently, preliminary codes were developed to capture key concepts and insights. These codes were refined and organized into broader, more comprehensive themes. From this iterative process, four principal themes emerged: (1) Collective health management through sharing practices; (2) navigating health information in a fragmented digital ecosystem; (3) digital PHI as a motivational tool for proactive health management; and (4) preference for traditional methods and perception of non-necessity. These themes highlighted how PHI is viewed and used within the context of personal health management among urban Indian participants. The thematic coding framework is detailed in Table 4, providing a structured overview of the identified themes and their subcategories.
3.4. Data Triangulation
To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature. This method enabled cross-verification and identification of patterns and discrepancies across data types, enhancing the validity of the results. For instance, survey and interview data revealed a lack of trust in wearable technologies due to perceived inaccuracies in advanced metrics. This notion was corroborated by existing literature that links health literacy and trust to digital health technologies (Bagalkot et al., 2020; Kreps, 2017). Our analysis of barriers to wearable adoption, such as a preference for traditional healthcare and the need for culturally tailored information, aligns with broader studies emphasizing context-specific approaches in non-Western settings (Ugargol and Bailey, 2018; Volpato et al., 2021). Further, we observed that participants often viewed wearables as supplementary rather than essential health tools for their PHI, reflecting a prevalent reactive approach to health management as shown by prior literature (Lee et al., 2022; Kang and Exworthy, 2022). By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings.
3.5. Positionality
The authors of this study acknowledge their positionality as researchers from the Indian cultural context. While some authors still live in India and bring an insider perspective, others live abroad with a background in distributed computing. This diverse positionality allowed for a comprehensive analysis considering local nuances and broader global trends in wearable technology use. The authors also recognize that their interpretations are influenced by their own experiences with using PHI and backgrounds, which may shape the analysis and conclusions drawn from the data. Throughout the research process, efforts were made to remain reflexive and mindful of potential biases, ensuring that the participants’ voices were accurately represented and that the analysis remained grounded in the data.
4. Findings
Using data triangulation from the various components of the study, we examined the perceptions, usage patterns, motivations, and barriers associated with PHI through an in-depth analysis. The thematic analysis revealed several key themes and sub-themes that offer valuable insights into how users interact with PHI, their trust in the technology, and their challenges/barriers to adoption.
4.1. Collective Health Management Through Sharing Practices
4.1.1. Health Seen as a Collective Responsibility
While major research has focused on individualistic healthcare models, the reflection of Indian practices indicates that health and wellbeing are shared/collective in nature. People often view the health of their family members as a shared responsibility, driving the need to keep on top of each other’s health statuses. For example, in our survey, over 60% (55/87) of participants were already tracking their health using a wearable, and in the interviews, about 45% (10/22) were gifted their tracking device to be used within families to track vital metrics, enabling a collaborative care model.
They viewed health management as a shared, collective process rather than an individual responsibility. They often rely on digital PHI not only to monitor their health but also to care for family members and close contacts. This shared approach fosters mutual support, accountability, and informed decision-making among trusted social circles. For instance, P8 disclosed that apart from getting a personal tracker, he also wanted to track his parents’ vitals to check on them even when living away.
”I also got a fitbit watch for my parents, and my parents wear them. So I can track heart rate and other metrics. So first, I bought for them, and then I got one for myself.” - P8 (M, 26)
Even among people who leave the country for work, their needs based on cultural values lead them to seek ways to track their parents’ health and support their health management remotely. This feeling was expressed through CP12’s personal experience of wanting to monitor his father’s health when he was living abroad a decade ago, but finding a lack of cost-effective means to do so.
”During my initial career days (living abroad), as my dad got surgery, I wanted to be able to monitor him regularly. But there was no cheap and reliable means to do so. So, I had to rely on relatives to inform me of his status. If there was an economical way to monitor my father continuously, it would have helped me a lot.” - CP12 (M, 40)
4.1.2. Understanding the Aspects of Sharing
Through the secondary interviews, co-design sessions on sharing preferences, and a Sharing Awareness Survey, we discovered that individuals are open to sharing their PHI with spouses and doctors without reservation, but are unaware of existing mechanisms to share. They had varied opinions on how they would like their data to be shared and used. Some couples (2/6) in the co-design session were open to their data being used by others as long as they were informed of the use case, while all the rest wanted control over not only sharing but also the storing and usage of the data. The participants’ interest in viewing all kinds of health information about themselves and their family members in a single place showcases a need for a unified system.
The level of openness with sharing decreased in the following order among the participants: Spouses Doctors Parents Siblings Friends Employer Insurance Companies. Some couples (2/6) also mentioned that they wanted to avoid sharing with employers and insurance companies if that was the option.
”So, I am open to sharing everything with my spouse. Maybe just an update on how I am doing with my parents maybe. I don’t think I’ll be sharing much with my siblings or friends except if there are any emergency alerts. I am also open to sharing everything with doctors as long as it makes sense for the doctor to be seeing those for the purpose of the visit.” - CP1 (F, 34)
When inquired about their knowledge and usage of any current sharing frameworks, most of the couples (5/6) were unaware of those already in the market through prominent brands (Apple Health and Samsung Health). This general lack of awareness and usage of such sharing frameworks was also reinforced through the outcome of the survey responses (94/114) from the health sharing feature survey. The aware couple said they were already using it (however sparingly) for things that can be tracked through their wearable. However, they indicated they are unhappy about the lack of control over what can be shared, how it can be shared, and when it can be shared. This observation hints that either the existing platforms are so rigid that users are unhappily using them as little as possible, or they are unaware of it because it is not popular in the urban Indian context. A couple even drew insights from their experience living abroad and the system they have seen being used.
”Sharing within spouses is OK, but what about the other stakeholders like employers or insurance companies? do we get a say on what data goes to them or how they use it? That is what I am looking for, if there is an application to seamless share my health information. See, for example, when we were in the UK, there was this app that shows all details concerning our GP, from appointments to medication. When we want to share the medication information with the chemist/pharmacist, we just need to hit share, and we can select what we want to share. That level of control would be very useful.” - CP6 (M, 30) & CP5 (F, 27)
Another insightful observation among participants interested in sharing was that people with children are interested in tracking and sharing PHI only when there is no overhead of pulling all things together by themselves. While the motivation to track using wearables differs individually (even within a couple), they all were motivated for the overall PHI for themselves and their children if they could do so using a single device.
”See, in this digital age, almost every data is online. Not only with each other but also with companies like Google or Apple (laughs). After this interview, I will probably get an Instagram reel on how couples should track data together (laughs). So, there is no need for stricter control within the couple on how/what/when to share. We currently use whatever means possible (mostly just WhatsApp each other on health documents and information), we are open to a dedicated platform only if we do not have to add everything, every time manually.” - CP7 (M, 33) & CP8 (F, 31)
After incorporating the feedback and inputs, when evaluating the rapid figma prototype, the participants liked the overall functionality of the interface and provided further suggestions. They felt that the prototype mimicked the discussion on their sharing preferences and appreciated the control over them. Some participants also suggested that further information or health tips could be added about the benefits of good sleep and (for women) steps to take when periods are late.
”Some family tracking and reward like a badge to the member who completes all goals, just to have encouraging competition in the family for better health would be nice.” - CP1 (F, 34)
”I like the daily, weekly, and monthly stats being shown for health features. But I would also like to have a graphical representation.” - (CP3, F, 33)
The responses from the Sharing Feature Awareness Survey also indicate a need for a reliable medium of sharing information among Android users and a better demonstration of capabilities among Apple/iPhone users.
”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage. If they can be stored all together in one place, that would be a helpful feature. ” - (F, 33)
”I thought sharing was about putting it out on social media platforms. I never knew about sharing features that send health updates to our family or health care provider.” - (M, 24)
Overall, participants’ motivations and needs indicate a need for a single dedicated application or a sharing mechanism where they can pool all kinds of health information about themselves and their trusted circles while maintaining full control over the preferences, such that they do not have additional cognitive load on managing it. While some prefer textual representation of this information, others are comfortable with graphical, while a very limited few are okay with numbers and medical jargon.
4.2. Navigating Health Information in a Fragmented Digital Ecosystem
4.2.1. Unreliable Tools and Systems
Users encounter a patchwork of digital health information, often leaving them uncertain about what to trust. While basic functionalities (like step counting and heart rate monitoring) are widely accepted, there is notable skepticism around advanced metrics. Participants frequently cross-check data with multiple devices, healthcare professionals, and even traditional methods to validate accuracy. This verification process underscores the challenge of navigating a fragmented digital health landscape.
”So I think, What is missing in current systems is having that authenticity or maybe accuracy? I don’t blindly believe them, like, OK, it’s saying something is wrong, so it must be, but I won’t ignore it either. If my tracker is giving some wrong data, I will check with another and see whether it’s me or it’s the sensor, but if 2 or 3 trackers have been giving you that data, I don’t think there is any need to doubt it. (laughs).” - (P4, F, 33)
Even the participants who have used various trackers throughout their evolution noted that the current capabilities of PHI systems are not on par with their current purpose or motivation. They emphasized the lack of reliable sources for health information or even the capability of systems to explain their health in understandable terms.
”So one of my first smartwatches was a Pebble smartwatch. It is very old. It started on Kickstarter (—) So that is how I got into it. (—) But more lately. It’s been to serve the purpose of. OK, how do I understand my body a little better? (—) So what it started with and what it is right now, it’s very different. The current systems cannot explain it well to me.” - P19 (M, 36)
4.2.2. Island of Ecosystems
Many of these PHI trackers have their own ecosystems that are specialized for a specific brand or product line, focused on boosting their unique functionalities to promote their line of products. This causes a lack of cross-functional capabilities. This was echoed during the interviews with doctors with different specialties, the common recurring theme was the existing use of specialized trackers for different chronic conditions like diabetes, digital pressure cuffs for hypertension, or even mood journaling for psychiatry. The doctors from endocrinology and medicine confirmed using existing specialized smart devices in practice, albeit their price factor limited it to a few patients. The high financial investment required for reliable devices further compounds these reservations for individuals as well.
”It’s kind of an investment, like it is at least for me. So buying a watch which which will cost 20,000/25,000 at the time of the release, it is kind of an expensive affair. So you cannot change these things with every update. ” - P3 (F, 33)
Wearables, an integral part of PHI for many users, had technical issues apart from their limitations of cross-functionality on what they could offer the user, causing further frustrations among participants. One such instance is inaccurate heart rate measurements needing recalibration during exercise. This observation shows that mechanisms such as comparison tools or validation indicators that help users assess the accuracy of their health data are currently lacking.
”It happens quite a lot, especially in the gym when I can feel that my heart is like pounding (laughs); like help!. OK, so I can feel my heart beating really fast. But when I see my watch, it will show readings like 75 or 76, which is kind of normal. (—) Then, then I have to again, again like reconfigure that (manually calibrate the reading). Then you put, then you press on that measure. So some kind of calibration is required.” - P3 (F, 33)
Participants noted the lack of reliable health information sources for them to verify their PHI and the reduced level of interoperability among available devices. They expressed that they do not have verified, reputable health information sources to compare their metrics and rely on doctor visits to make sense of them.
”I do not understand whether the heart rate is good for me or not, the tracker does not tell me that. When I had the time to go to my doctor that time also I was made to understand that it is in between 72 and 80, which was fine according to the doctor. Presently also, it is nearly like that. When I walk, it shows 78 to 82; when I am not, it shows something about 68 to 72. So, I am guessing it is OK, even though I am not 100% sure.” - (P10, M, 62)
Overall, the observations with participants’ struggles to make sense of the data indicate that the fragmented ecosystems are complex and lack inter-compatibility to use them all cohesively for PHI. Simplifying the complexity of digital health data through intuitive design and user-friendly explanations that boost both digital and health literacy could aid in their journey better, as noted by a few participants.
”So if you get a chance to kind of have that information to help you navigate through that (PHI), it will be great. Because, at the end of the day, being healthier is everyone’s priority. The lack of such information or even the means to get them is stopping people (from proactive health management), with other priorities getting ahead.” - (P9, F, 29)
4.3. Digital PHI as a Motivational Tool for Proactive Health Management
Digital PHI catalyzes behavior change by transforming passive data collection into proactive health management. Users appreciate features that provide real-time feedback, set personalized goals, and track progress over time. These tools encourage people to adopt healthier lifestyles, offering actionable insights and reminders. However, the motivational impact is closely tied to individuals’ priorities and preferences. For example, some participants were mindful of their lifestyle, like P16 (F, 26) and P7 (M,32), who track their steps and maintain an active lifestyle during sedentary workdays.
”Since I have a desk job, so I have to make sure that I’m doing at least, like so far that I come across the Internet, they do mention that you should be at least walking 5 kilometers if you are sitting all day and working. So that is one thing I keep in mind. That’s what makes me keep a check on my step count and how much I’m walking.” - (P16, F, 26)
”So I was walking very fast. Then it (the tracker) told me you are walking outdoors; you have walked so much distance. So then, I wish to walk more.” - (P7, M, 32)
Integrating reminders, gamified challenges, and progress-tracking features that motivate users to stay engaged with their health data was among the recommendations from participants in the interviews and rapid prototyping. They considered them resourceful motivators, a means to make healthy habits, or a means for a more personalized health management tool by delivering tailored advice or actionable insights that resonate with their individual and family health goals.
”So the (health) reminder thing, right, it can be a good guilt-tripping system so. (—) more like a motivator (to work out more). This would be a good, good thing to see, like a progress of myself.” - (P17, M, 26)
Participants were eager to view summarized stats over day, week, and month. However, they also wished for them to be presented alternately to increase their engagement. Participants noted that being handy could also be another reason for adopting digital PHI as a reliable tool. This finding indicates the potential of digital PHI as a motivational tool if different presentation modalities are provided.
4.4. Preference for Traditional Methods and Perception of Non-necessity
4.4.1. Traditional and Cultural Influence
Some participants remain cautious about fully adopting digital PHI, favoring traditional health management methods. Concerns about privacy, technological complexity, and physical discomfort (e.g., frequent charging needs and skin irritation) contribute to this resistance. Some users (5) trust their bodily awareness and conventional healthcare practices, finding digital tracking unnecessary or overwhelming.
”I do not prefer to constantly check my miband. I only use it when I run or perform any exercise. I do not find the need to check the history or track, as if something is wrong, I could feel it in my body.” - P21(M,27)
Similarly, P20 (M, 29) emphasized his belief that his body would alert him if something were wrong, reducing the perceived need for a smartwatch and confidence in his healthier lifestyle.
”Whenever you know I’m like whenever I’m sick, (laughs) maybe at the time, something is wrong and I don’t, as you said (tracking health metrics regularly) maybe I’m not much bothered about, the specifics. (—) If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad. So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
Some participants even recognized and reasoned about the disjointedness between the customs of their parents’ generation and the current one, indicating a need for better integrating digital PHI with traditional cultural practices to promote active, healthy living.
”If you think about it, our lifestyles has been completely different from our parents. These (PHI) trackers only work for us. Designers are completely ignoring the population that still leads a more self-reliant lifestyle. If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
While participants who seek a more body-over-technology approach exist, they also make recommendations in ways that can improve for a wider audience. This observation shows that the adoption is dictated by the value it offers, which does not meet their needs, and indicates room for improvement.
4.4.2. Barriers or Challenges Causing Digital PHI Resistance
Participants considered an assurance of precision in both basic and advanced features a necessity to maintain trust and encourage ongoing proactive engagement. They compared their childhood experience with traditional measuring devices as a frame of reference for building trust in digital PHI systems, indicating a need to get used to such digital systems.
”Maybe we are just used to a different kind of devices. For example, growing up, for the blood pressure machine (makes hand motions of the pumping used in a blood pressure cuff) (laughs), that thing, like you see something happening. (—) I think it’s one of the perceptive Issues. And also because I’m not used it a lot. I don’t have the perception, and its that.” - (P5, F, 29)
Comfort and lifestyle fit have also emerged as critical factors influencing the continued use of smart wearables as care devices, causing the aspects of using a wearable to become a barrier for a few. For example, P9 (F, 29), who received a smartwatch as a gift, eventually found the device’s frequent charging needs and complexities overwhelming, not integrating seamlessly into her life.
”So, so one, once I basically invested in this watch. I mean it was all fancy and all of it (—) What I realized is that this was too much of a task for me. The most painful thing is to kind of charge the device. (—) I wanted to do something to kind of keep my health in check, but it got too much for me.” - P9 (F,29)
For others, the physical discomfort or rashes due to wearable material can detract from their overall utility, and others prefer different modalities such as rings, bands, or other wearables to a watch so they can wear their traditional watches, which they enjoy wearing.
”For me, specifically, I do not prefer to wear it continuously because I get rashes from the strap if I wear it for too long regardless of the kind of strap.” - P22 (F,27)
”I’m also trying to go away from a watch format and more a band or a ring or something else of that sort. (—) Mostly because I also wanna like continue wearing my other watches.” - (P19, M, 36)
Overall, the findings indicate that digital PHI can offer options that allow for both proactive tracking and reactive monitoring, acknowledging users’ comfort with conventional health cues would be welcome among users. They also recommend cost-effective solutions to ease the transition from traditional to digital health management.
5. Discussion
Our findings shed light on how urban Indians adopt, use, and perceive health wearables for PHI in India, emphasizing the cultural, societal, and technological factors influencing user behavior. While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we uncover a broader gap: how these same technologies are—or could be—woven into routine health management practices. By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape PHI adoption. Below, we discuss three key themes—navigating a fragmented ecosystem, wearables as care devices, and culturally sensitive engagement —and offer design and policy recommendations to support a more culturally attuned PHI future.
5.1. Navigating a Fragmented Digital Health Ecosystem: Technology & Policy
India’s digital health landscape is marked by fragmentation and the absence of centralized, reliable information—conditions that contrast sharply with the integrated systems typical in Western contexts. This fragmentation hampers PHI usability, trust, and effectiveness.
Participants reported difficulty navigating multiple disjointed platforms, often defaulting to traditional health practices. Choosing a tracker was perceived as a significant investment, complicated by non-compatible ecosystems—a clear barrier (section 4.2). While PHI systems evolve to support personal health management, the IT infrastructure backing them remains outdated and fragmented (Jones et al., 2012). Proprietary ecosystems hinder interoperability, complicating and increasing the cost of decision-making. For example, a Samsung smartwatch user with an iPhone or non-Samsung Android device cannot access all features or rely on data accuracy, illustrating how such silos generate user confusion and disengagement. This calls for design strategies that reduce the cognitive load of choice and simplify use. As seen in our prototype evaluation, users need flexible interaction modes and wearable form factors tailored to their comfort levels (section 4.4.2). While some of these challenges can be overcome through software/interaction design, the rest require policy-level changes that enforce manufacturers to comply with an open data framework that ensures accessibility of health data regardless of the medium or proprietary ecosystem requirements of individual brands/companies.
Furthermore, the fragmented digital ecosystem and the accompanying mistrust raise ethical and privacy concerns. This aspect of PHI must be balanced with the need for integrated, actionable health information. While countries worldwide have policies regarding collecting and using health information (for example, HIPAA, PHIPA, GDPR, and APPs), India still uses the IT Act of 2000 even for medical and health information. The IT Act 2000 provides basic data protection by requiring data owners to be informed about its use, but it is not tailored to healthcare. The National Digital Health Blueprint (NDHB) recommends ISO 22600:2014 Health Informatics, but implementing these standards may be challenging for many healthcare providers in India, and ensuring compliance at the provider level could also be challenging for the government. Although similar healthcare-specific legislation was proposed in 2017 and 2022, it has not been implemented. While the Government of India has now published the “Digital Personal Data Protection Act, 2023” (DPDP 2023) covering all personal data (including health data), it remains unenforced and is not robust enough to meet the requirements posed by Indian users’ socially reflective shared caring models.
We propose a unified, open framework that aggregates PHI, doctor visits, lab reports, medications, and more into a single interface. This interface should support all smart wearables—focused on Android, which dominates the Indian market (94%)—and employs accessible language to reduce learning burdens and foster proactive health behaviors.
5.2. Wearables as Care Devices: Collective Care & Engagement Strategies
Our participants overwhelmingly described health wearables not merely as personal gadgets but as instruments of family‐centered care (findings 4.1.1). Unlike the individualist framing common in Western studies (Chmielewski and Meyer, 2024), Indian users envision sharing health metrics/analysis with parents, spouses, or children, turning data streams into collaborative decision‐making tools. We observed that users readily assume caregiver roles—monitoring a parent’s daily steps or a child’s sleep patterns —and express enthusiasm for co‐designed interfaces that let family dyads explore and interpret data together (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023).
To drive proactive engagement, gamification of PHI systems must capture activities familiar to Indian lifestyles—such as traditional labor and household tasks—rather than only Western‐style workouts to motivate them better (findings 4.1.2 & 4.3). At the same time, designs should include contextual alerts (for example, flagging sustained deviations in resting heart rate) that prompt timely medical check‐ins, moving users from reactive responses toward ongoing health management. Prior studies on intergenerational sharing of health data (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024a), women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a), and children-specific care for chronic ailments (Cha et al., 2024), reinforce that broadening the scope of shareable metrics can foster more comprehensive, preventive care—particularly where caregiving is a shared, familial endeavor.
Privacy concerns were minimal among our participants (finding 4.1.2), indicating openness to data exchange within trusted circles. This suggests potential for paired or multi‐user wearable configurations, provided that the system offers clear, in-app explanations of how data are collected, processed, and used. Policymakers can foster privacy-conscious behaviors by enforcing clear regulations around data protection, specifically in PHI wearables. As people adopt a more proactive approach, wearables’ role in caregiving could be significantly enhanced by providing continuous insights that allow early intervention before critical incidents. The ability to share comprehensive health data with family members could also strengthen the perceived value of these systems, encouraging consistent use and deeper integration into daily routines. Designers and developers must therefore prioritize features that integrate into users’ lives, facilitating comprehensive health monitoring and sharing. Such usability and transparency, coupled with policy measures mandating robust PHI protections, will be essential to transforming wearables from novelty gadgets into indispensable tools for early intervention and collaborative health management. Our observations showcase that the ”Personal” in PHI involves more than the individual in the Indian context. These family-centered insights directly inform our unified PHI platform vision in the next section.
5.3. Culturally Sensitive Engagement: Norms, Literacy, Misinformation, & Trust
In India, paper-and-analog practices remain deeply trusted, but users embrace digital PHI when it feels like an extension, not a replacement. Participants, for example, described “smart” digital pressure cuffs that look and behave like traditional sphygmomanometers yet auto-sync readings to a unified PHI dashboard (findings 4.4.2). We frame this as a modular hybrid PHI paradigm, where users toggle between digital and analog interfaces based on context and comfort. Such an approach acknowledges India’s healthcare realities and offers a pragmatic, modular pathway for digital transformation—one that integrates seamlessly with generic platforms.
In a society where talking openly about intimate or sexual health is still taboo, cultural norms around sharing vary by metric: intimate health data (e.g., sexual health) was shared only with spouses, whereas vitals like step counts traveled freely among parents, children, or extended family depending on trust (findings 4.1.1). Parents even leaned on devices recommended by their children, illustrating “derived trust.” To embrace these differences, we argue for a culturally informed data-governance model with user-controlled sharing settings and transparent, in-app analytics—so families can collaborate on care without over- or under-sharing. Future policy frameworks should incorporate these principles to foster an ecosystem where digital health innovations are ethically grounded and culturally resonant.
Low health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) further compound these challenges and impact adoption. Participants distrusted novel features when metrics seemed “off” or unfamiliar/unevaluated (section 4.2.2), and authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently. This disparity highlights the need for authentic, reliable, vetted, user-friendly health information platforms so Indians can access trustworthy guidance without skepticism. And since the pandemic has driven more health-conscious behaviors, implementing the long-overdue 2005 school-health curriculum recommendation (MoHFW, 2005) could foster health-conscious behavior from a young age (Kann et al., 2007; Pulimeno et al., 2020; Auld et al., 2020; Benes and Alperin, 2022; Sulz et al., 2024).
Trust hinges on data accuracy and medical validation. Wearables often function as “black boxes,” limiting lay users’ comprehension (Shin et al., 2019; Zhang and Shahriar, 2020). Sustained adoption, therefore, requires medical validation and clinical integration. Devices should earn formal endorsements, sync with electronic health records, and surface sensor-accuracy indicators (e.g., confidence intervals around SpO2 or heart-rate) to bolster user confidence (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023). Integrating wearables with formal healthcare systems will further bolster their credibility and sustained use.
These collective/shared care insights directly inform our integrated PHI platform that consolidates disparate data, embeds user-controlled sharing and transparent analytics, employs inclusive design methods (e.g., family workshops), and links seamlessly with clinical systems—so that as health literacy grows, users can critically weigh benefits against risks (Mackert et al., 2016; Meyers et al., 2020).
5.4. Future Work
While there are novel recommendations on personal, public, and cultural aspects of PHI adoption, the authors acknowledge several limitations in this study. First, the cross-sectional design captures user behavior and perceptions at a single point in time, limiting the ability to observe how these factors may evolve. Additionally, the focus on urban, middle- and upper-middle-class participants restricts the generalizability of the findings to rural and lower-income populations, where different adoption barriers, such as different lifestyle practices, natural remedies, and many more, may exist. Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their PHI usage, utility, and adoption.
Building upon our observations and the strategic direction outlined, we propose the design and development of a proof-of-concept application aimed at addressing general health, wellbeing, and medical inquiries. This system will also assist users in interpreting and contextualizing their personal health data aggregated from diverse sources, presenting the information in a format that is both comprehensible and actionable. Leveraging advancements in artificial intelligence, particularly finetuned open-source large language models (LLMs), the application will integrate data from clinical documents (e.g., physician reports, laboratory results, and medication records) alongside information from wearable health devices. This multimodal data synthesis will enable the generation of personalized, holistic insights, thereby enhancing user engagement and health awareness. The interface will employ a conversational, chatbot-style design to facilitate intuitive interaction, mediate family sharing norms, and surface literacy-appropriate explanations. Beyond helping users understand their personal health information (PHI), the system will also provide access to evidence-based health, medical, and wellness content. By sourcing validated information from reputable sources, the application seeks to address the pervasive issue of health misinformation and contribute to the improvement of public health literacy.
6. Conclusion
This study comprehensively explores how Indian users integrate PHI wearables for health management. Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of PHI in India. By doing so, we emphasize the influence of sociocultural factors, usability challenges, and trust issues in their consistent/long-term use of PHI for health management. While PHI and specifically wearables hold significant potential to shift personal health management from a reactive to a proactive model, their integration is hindered by cultural norms, limited health literacy, and skepticism around data accuracy. To address these challenges, we propose a dual focus on improving health literacy and building trust through transparent communication, medical validation, and user-centered design.
Two essential contributions of this research are identifying critical areas where designers and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants. The critical areas include expanding PHI functionalities to align with family-based caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information. As shown by the evaluated prototype, the HCI community, technology developers, and policymakers must work together to standardize core functionalities and integrate PHI into formal healthcare systems. By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive health management, supporting both individual users and familial care structures.
This work offers a framework for future research and industry efforts, reinforcing the transformative potential of PHI within diverse cultural contexts. By addressing technological and sociocultural barriers, PHI can aid in preventive healthcare, fostering long-term engagement and broader adoption across India and beyond.
Acknowledgements.
We would like to thank the student Harsh Pala for his involvement in the development of Figma prototypes used for evaluating an interactive interface of the sharing preferences identified by the card sorting activity.References
- (1)
- Alrowaily (2024) Majed Abdullah Alrowaily. 2024. Smart Emergency Alerting System: A Machine Learning Approach. In Proceedings of the 2024 2nd International Conference on Communications, Computing and Artificial Intelligence (CCCAI ’24). Association for Computing Machinery, New York, NY, USA, 8–15. https://doi.org/10.1145/3676581.3676583
- Amato et al. (2022) Flora Amato, Walter Balzano, and Giovanni Cozzolino. 2022. Design of a Wearable Healthcare Emergency Detection Device for Elder Persons. Applied Sciences 12, 5 (Jan. 2022), 2345. https://doi.org/10.3390/app12052345
- Anderson and Durstine (2019) Elizabeth Anderson and J. Larry Durstine. 2019. Physical Activity, Exercise, and Chronic Diseases: A Brief Review. Sports Medicine and Health Science 1, 1 (Dec. 2019), 3–10. https://doi.org/10.1016/j.smhs.2019.08.006
- Ankrah et al. (2023) Elizabeth A. Ankrah, Franceli L. Cibrian, Lucas M. Silva, Arya Tavakoulnia, Jesus A. Beltran, Sabrina E.b. Schuck, Kimberley D. Lakes, and Gillian R. Hayes. 2023. Me, My Health, and My Watch: How Children with ADHD Understand Smartwatch Health Data. ACM Trans. Comput.-Hum. Interact. 30, 4 (Sept. 2023), 59:1–59:25. https://doi.org/10.1145/3577008
- Auld et al. (2020) M. Elaine Auld, Marin P. Allen, Cicily Hampton, J. Henry Montes, Cherylee Sherry, Angela D. Mickalide, Robert A. Logan, Wilma Alvarado-Little, and Kim Parson. 2020. Health Literacy and Health Education in Schools: Collaboration for Action. NAM Perspectives 2020 (July 2020), 10.31478/202007b. https://doi.org/10.31478/202007b
- Bagalkot et al. (2020) Naveen Bagalkot, Nervo Verdezoto, Anushri Ghode, Shipra Purohit, Lakshmi Murthy, Nicola Mackintosh, and Paula Griffiths. 2020. Beyond Health Literacy: Navigating Boundaries and Relationships During High-risk Pregnancies: Challenges and Opportunities for Digital Health in North-West India. In Proceedings of the 11th Nordic Conference on Human-Computer Interaction: Shaping Experiences, Shaping Society (NordiCHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3419249.3420126
- Bastico et al. (2022) Matteo Bastico, Verónica Ruiz Bejerano, and Alberto Belmonte-Hernández. 2022. Simultaneous Real-Time Human Fall Detection and Reidentification Based on Multisensors Data. In Proceedings of the 15th International Conference on PErvasive Technologies Related to Assistive Environments (PETRA ’22). Association for Computing Machinery, New York, NY, USA, 365–370. https://doi.org/10.1145/3529190.3534728
- Bat-Erdene and Saver (2021) Bat-Orgil Bat-Erdene and Jeffrey L. Saver. 2021. Automatic Acute Stroke Symptom Detection and Emergency Medical Systems Alerting by Mobile Health Technologies: A Review. Journal of Stroke and Cerebrovascular Diseases 30, 7 (July 2021), 105826. https://doi.org/10.1016/j.jstrokecerebrovasdis.2021.105826
- Benes and Alperin (2022) Sarah Benes and Holly Alperin. 2022. The Essentials of Teaching Health Education: Curriculum, Instruction, and Assessment. Human Kinetics, United States.
- Bhat et al. (2023) Karthik S. Bhat, Amanda K. Hall, Tiffany Kuo, and Neha Kumar. 2023. ”We Are Half-Doctors”: Family Caregivers as Boundary Actors in Chronic Disease Management. Proc. ACM Hum.-Comput. Interact. 7, CSCW1 (April 2023), 111:1–111:29. https://doi.org/10.1145/3579545
- Bhat and Kumar (2020) Karthik S. Bhat and Neha Kumar. 2020. Sociocultural Dimensions of Tracking Health and Taking Care. Proc. ACM Hum.-Comput. Interact. 4, CSCW2 (Oct. 2020), 129:1–129:24. https://doi.org/10.1145/3415200
- Binda et al. (2017) Jomara Binda, Hyehyun Park, John M. Carroll, Natalie Cope, Chien Wen (Tina) Yuan, and Eun Kyoung Choe. 2017. Intergenerational Sharing of Health Data among Family Members. In Proceedings of the 11th EAI International Conference on Pervasive Computing Technologies for Healthcare (PervasiveHealth ’17). Association for Computing Machinery, New York, NY, USA, 468–471. https://doi.org/10.1145/3154862.3154895
- Blandford (2019) Ann Blandford. 2019. HCI for Health and Wellbeing: Challenges and Opportunities. International Journal of Human-Computer Studies 131 (Nov. 2019), 41–51. https://doi.org/10.1016/j.ijhcs.2019.06.007
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using Thematic Analysis in Psychology. Qualitative Research in Psychology 3, 2 (Jan. 2006), 77–101. https://doi.org/10.1191/1478088706qp063oa
- Bronfenbrenner (1974) Urie Bronfenbrenner. 1974. Developmental Research, Public Policy, and the Ecology of Childhood. Child Development 45, 1 (1974), 1–5. https://doi.org/10.2307/1127743
- Brun and Häkkilä (2022) Damien Brun and Jonna Häkkilä. 2022. Exploration of Electrochromics for Calm Reminder: A Customizable Bracelet. In Proceedings of the 33rd Australian Conference on Human-Computer Interaction (OzCHI ’21). Association for Computing Machinery, New York, NY, USA, 269–272. https://doi.org/10.1145/3520495.3520509
- Cagiltay and Mutlu (2024) Bengisu Cagiltay and Bilge Mutlu. 2024. Toward Family-Robot Interactions: A Family-Centered Framework in HRI. In Proceedings of the 2024 ACM/IEEE International Conference on Human-Robot Interaction (HRI ’24). Association for Computing Machinery, New York, NY, USA, 76–85. https://doi.org/10.1145/3610977.3634976
- Carter et al. (2014) Nancy Carter, Denise Bryant-Lukosius, Alba DiCenso, Jennifer Blythe, and Alan J. Neville. 2014. The Use of Triangulation in Qualitative Research. Oncology Nursing Forum 41, 5 (Sept. 2014), 545–547. https://doi.org/10.1188/14.ONF.545-547
- Cha et al. (2024) Yoon Jeong Cha, Yasemin Gunal, Alice Wou, Joyce Lee, Mark W Newman, and Sun Young Park. 2024. Shared Responsibility in Collaborative Tracking for Children with Type 1 Diabetes and Their Parents. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–20. https://doi.org/10.1145/3613904.3642344
- Chandrasekaran et al. (2020) Ranganathan Chandrasekaran, Vipanchi Katthula, and Evangelos Moustakas. 2020. Patterns of Use and Key Predictors for the Use of Wearable Health Care Devices by US Adults: Insights from a National Survey. Journal of Medical Internet Research 22, 10 (Oct. 2020), e22443. https://doi.org/10.2196/22443
- Cherian et al. (2024) Josh Cherian, Samantha Ray, Thomas Mernar, Paul Taele, Helen Mach, Jung In Koh, Ping Ye, and Tracy Hammond. 2024. A Step Toward Better Care: Understanding What Caregivers and Residents in Assisted Living Facilities Value in Health Monitoring Systems. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 13:1–13:29. https://doi.org/10.1145/3637290
- Chmielewski and Meyer (2024) Marta Chmielewski and Matthew J. Meyer. 2024. Discussions With End Users to Inform the Vision for a Shared Care Record in Ontario: Qualitative Interview Study. Online Journal of Public Health Informatics 16, 1 (Feb. 2024), e51231. https://doi.org/10.2196/51231
- Costa Figueiredo and Chen (2021) Mayara Costa Figueiredo and Yunan Chen. 2021. Health Data in Fertility Care: An Ecological Perspective. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3411764.3445189
- Creaser et al. (2022) Amy V. Creaser, Jennifer Hall, Silvia Costa, Daniel D. Bingham, and Stacy A. Clemes. 2022. Exploring Families’ Acceptance of Wearable Activity Trackers: A Mixed-Methods Study. International Journal of Environmental Research and Public Health 19, 6 (March 2022), 3472. https://doi.org/10.3390/ijerph19063472
- Delahoz and Labrador (2014) Yueng Santiago Delahoz and Miguel Angel Labrador. 2014. Survey on Fall Detection and Fall Prevention Using Wearable and External Sensors. Sensors 14, 10 (Oct. 2014), 19806–19842. https://doi.org/10.3390/s141019806
- Egede et al. (2021) Joy Egede, Maria J. Galvez Trigo, Adrian Hazzard, Martin Porcheron, Edgar Bodiaj, Joel E. Fischer, Chris Greenhalgh, and Michel Valstar. 2021. Designing an Adaptive Embodied Conversational Agent for Health Literacy: A User Study. In Proceedings of the 21st ACM International Conference on Intelligent Virtual Agents (IVA ’21). Association for Computing Machinery, New York, NY, USA, 112–119. https://doi.org/10.1145/3472306.3478350
- Epstein et al. (2015) Daniel A. Epstein, An Ping, James Fogarty, and Sean A. Munson. 2015. A Lived Informatics Model of Personal Informatics. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 731–742. https://doi.org/10.1145/2750858.2804250
- Evans et al. (2022) Hayley Irene Evans, Catherine R Deeter, Jiawei Zhou, Kimberly Do, Andrew M Sherrill, and Rosa I. Arriaga. 2022. Perspectives on Integrating Trusted Other Feedback in Therapy for Veterans with PTSD. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3491102.3517513
- Fritz et al. (2014) Thomas Fritz, Elaine M. Huang, Gail C. Murphy, and Thomas Zimmermann. 2014. Persuasive Technology in the Real World: A Study of Long-Term Use of Activity Sensing Devices for Fitness. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’14). Association for Computing Machinery, New York, NY, USA, 487–496. https://doi.org/10.1145/2556288.2557383
- Ghosh and Das (2024) Manthan Ghosh and Deepa Das. 2024. Voice-Activated SOS: An AI-Enabled Wearable Device. In Impact of AI on Advancing Women’s Safety. IGI Global, Pennsylvania, USA, 251–277. https://doi.org/10.4018/979-8-3693-2679-4.ch016
- Gjoreski et al. (2016) Hristijan Gjoreski, Jani Bizjak, and Matja Gams. 2016. Using Smartwatch as Telecare and Fall Detection Device. In 2016 12th International Conference on Intelligent Environments (IE). IEEE, London, UK, 242–245. https://doi.org/10.1109/IE.2016.55
- Gooch et al. (2021) Daniel Gooch, Blaine A. Price, Anna Klis-Davies, and Julie Webb. 2021. A Design Exploration of Health-Related Community Displays. Proc. ACM Hum.-Comput. Interact. 5, CSCW1 (April 2021), 85:1–85:22. https://doi.org/10.1145/3449159
- Grimme et al. (2024) Sophie Grimme, Susanna Marie Spoerl, Susanne Boll, and Marion Koelle. 2024. My Data, My Choice, My Insights: Women’s Requirements When Collecting, Interpreting and Sharing Their Personal Health Data. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–18. https://doi.org/10.1145/3613904.3642851
- Gustafsson et al. (2022) Lena-Karin Gustafsson, Ildikó Asztalos Morell, Carl Johansson, and Santa De. 2022. Informal Caregiving from the Perspectives of Older People Living Alone in India. International Journal of Older People Nursing 17, 6 (Nov. 2022), e12468. https://doi.org/10.1111/opn.12468
- Haque et al. (2024) MD Romael Haque, Zeno Franco, Praveen Madiraju, Natalie D Baker, Sheikh Iqbal Ahamed, Otis Winstead, Robert Curry, and Sabirat Rubya. 2024. ”Butt Call Me Once You Get a Chance to Chat 🙂” : Designing Persuasive Reminders for Veterans to Facilitate Peer-Mentor Support. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3613904.3642962
- Hoque and Bao (2015) Md. Rakibul Hoque and Yukun Bao. 2015. Cultural Influence on Adoption and Use of E-Health: Evidence in Bangladesh. Telemedicine and e-Health 21, 10 (Oct. 2015), 845–851. https://doi.org/10.1089/tmj.2014.0128
- Huang et al. (2022) Yue Huang, Borke Obada-Obieh, Elissa M. Redmiles, Satya Lokam, and Konstantin Beznosov. 2022. COVID-19 Information-Tracking Solutions: A Qualitative Investigation of the Factors Influencing People’s Adoption Intention. In Proceedings of the 2022 Conference on Human Information Interaction and Retrieval (CHIIR ’22). Association for Computing Machinery, New York, NY, USA, 12–24. https://doi.org/10.1145/3498366.3505756
- Huhn et al. (2022) Sophie Huhn, Miriam Axt, Hanns-Christian Gunga, Martina Anna Maggioni, Stephen Munga, David Obor, Ali Sié, Valentin Boudo, Aditi Bunker, Rainer Sauerborn, Till Bärnighausen, and Sandra Barteit. 2022. The Impact of Wearable Technologies in Health Research: Scoping Review. JMIR mHealth and uHealth 10, 1 (Jan. 2022), e34384. https://doi.org/10.2196/34384
- Hwang et al. (2024) Angel Hsing-Chi Hwang, Dan Adler, Meir Friedenberg, and Qian Yang. 2024. Societal-Scale Human-AI Interaction Design? How Hospitals and Companies Are Integrating Pervasive Sensing into Mental Healthcare. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3613904.3642793
- Jones et al. (2012) Spencer Jones, Paul Heaton, Robert Rudin, and Eric Schneider. 2012. Unraveling the IT Productivity Paradox - Lessons for Health Care. The New England journal of medicine 366 (June 2012), 2243–5. https://doi.org/10.1056/NEJMp1204980
- Jones et al. (2021) Simon L. Jones, William Hue, Ryan M. Kelly, Rosemarie Barnett, Violet Henderson, and Raj Sengupta. 2021. Determinants of Longitudinal Adherence in Smartphone-Based Self-Tracking for Chronic Health Conditions: Evidence from Axial Spondyloarthritis. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 5, 1 (March 2021), 16:1–16:24. https://doi.org/10.1145/3448093
- Kabir and Wiese (2023) Kazi Sinthia Kabir and Jason Wiese. 2023. A Meta-Synthesis of the Barriers and Facilitators for Personal Informatics Systems. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 3 (Sept. 2023), 103:1–103:35. https://doi.org/10.1145/3610893
- Kang and Exworthy (2022) Harjeevan Singh Kang and Mark Exworthy. 2022. Wearing the Future—Wearables to Empower Users to Take Greater Responsibility for Their Health and Care: Scoping Review. JMIR mHealth and uHealth 10, 7 (July 2022), e35684. https://doi.org/10.2196/35684
- Kann et al. (2007) Laura Kann, Susan K. Telljohann, and Susan F. Wooley. 2007. Health Education: Results From the School Health Policies and Programs Study 2006. Journal of School Health 77, 8 (2007), 408–434. https://doi.org/10.1111/j.1746-1561.2007.00228.x
- Kapania et al. (2022) Shivani Kapania, Oliver Siy, Gabe Clapper, Azhagu Meena SP, and Nithya Sambasivan. 2022. ”Because AI Is 100% Right and Safe”: User Attitudes and Sources of AI Authority in India. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–18. https://doi.org/10.1145/3491102.3517533
- Karnatak et al. (2023) Nimisha Karnatak, Brooke Loughrin, Tiffany Amy Kuo, Odeline Mateu-Silvernail, Indrani Medhi Thies, William Thies, and Mohit Jain. 2023. “Is It Even Giving the Correct Reading or Not?”: How Trust and Relationships Mediate Blood Pressure Management in India. ACM Trans. Comput.-Hum. Interact. 30, 6 (Sept. 2023), 90:1–90:27. https://doi.org/10.1145/3609327
- Karusala et al. (2023a) Naveena Karusala, Victoria G, Shirley Yan, and Richard Anderson. 2023a. Unsettling Care Infrastructures: From the Individual to the Structural in a Digital Maternal and Child Health Intervention. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581553
- Karusala et al. (2023b) Naveena Karusala, Shirley Yan, Nupoor Rajkumar, Victoria G, and Richard Anderson. 2023b. Speculating with Care: Worker-centered Perspectives on Scale in a Chat-based Health Information Service. Proc. ACM Hum.-Comput. Interact. 7, CSCW2 (Oct. 2023), 361:1–361:26. https://doi.org/10.1145/3610210
- Kemp et al. (2021) Emma Kemp, Joshua Trigg, Lisa Beatty, Chris Christensen, Haryana M. Dhillon, Anthony Maeder, Patricia A. H. Williams, and Bogda Koczwara. 2021. Health Literacy, Digital Health Literacy and the Implementation of Digital Health Technologies in Cancer Care: The Need for a Strategic Approach. Health Promotion Journal of Australia 32, S1 (2021), 104–114. https://doi.org/10.1002/hpja.387
- Kim et al. (2022) Jeong-Kyun Kim, Da-Som Oh, Kangbok Lee, and Sang Gi Hong. 2022. Fall Detection Based on Interpretation of Important Features with Wrist-Wearable Sensors. In Proceedings of the 28th Annual International Conference on Mobile Computing And Networking (MobiCom ’22). Association for Computing Machinery, New York, NY, USA, 823–825. https://doi.org/10.1145/3495243.3558250
- Klasnja et al. (2011) Predrag Klasnja, Sunny Consolvo, and Wanda Pratt. 2011. How to Evaluate Technologies for Health Behavior Change in HCI Research. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’11). Association for Computing Machinery, New York, NY, USA, 3063–3072. https://doi.org/10.1145/1978942.1979396
- Kreps (2017) Gary L. Kreps. 2017. The Relevance of Health Literacy to mHealth. Information Services & Use 37, 2 (Jan. 2017), 123–130. https://doi.org/10.3233/ISU-170828
- Kruzan et al. (2023) Kaylee Payne Kruzan, Ada Ng, Colleen Stiles-Shields, Emily G Lattie, David C. Mohr, and Madhu Reddy. 2023. The Perceived Utility of Smartphone and Wearable Sensor Data in Digital Self-tracking Technologies for Mental Health. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581209
- Kyytsönen et al. (2023) Maiju Kyytsönen, Tuulikki Vehko, Heidi Anttila, and Jonna Ikonen. 2023. Factors Associated with Use of Wearable Technology to Support Activity, Well-Being, or a Healthy Lifestyle in the Adult Population and among Older Adults. PLOS Digital Health 2, 5 (May 2023), e0000245. https://doi.org/10.1371/journal.pdig.0000245
- Lazar et al. (2015) Amanda Lazar, Christian Koehler, Theresa Jean Tanenbaum, and David H. Nguyen. 2015. Why We Use and Abandon Smart Devices. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 635–646. https://doi.org/10.1145/2750858.2804288
- Lee et al. (2024a) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee. 2024a. FamilyScope: Visualizing Affective Aspects of Family Social Interactions Using Passive Sensor Data. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 57:1–57:27. https://doi.org/10.1145/3637334
- Lee et al. (2024b) Jarrett G.W. Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe. 2024b. Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2 (May 2024), 51:1–51:21. https://doi.org/10.1145/3659585
- Lee et al. (2022) Wan Ling Lee, Zi Jing Lim, Li Yoong Tang, Nor Aziyan Yahya, Kasturi Dewi Varathan, and Salizar Mohamed Ludin. 2022. Patients’ Technology Readiness and eHealth Literacy: Implications for Adoption and Deployment of eHealth in the COVID-19 Era and Beyond. CIN: Computers, Informatics, Nursing 40, 4 (April 2022), 244. https://doi.org/10.1097/CIN.0000000000000854
- Lieber et al. (2020) Judith Lieber, Lynda Clarke, Ian M. Timæus, Poppy Alice Carson Mallinson, and Sanjay Kinra. 2020. Changing Family Structures and Self-Rated Health of India’s Older Population (1995-96 to 2014). SSM - Population Health 11 (Aug. 2020), 15. https://doi.org/10.1016/j.ssmph.2020.100572
- Lu et al. (2024) Yuqian Lu, Zihan Zhou, and Xuelian Li. 2024. Study on the Relationship between Technology, Data and Body Isomorphism—The Example of Smartwatches. Telematics and Informatics Reports 13 (March 2024), 100118. https://doi.org/10.1016/j.teler.2024.100118
- Luz and Masoodian (2024) Saturnino Luz and Masood Masoodian. 2024. Managing Personal Health and Well-being by Integrating Data Streams Using Interactive Visualizations. In Proceedings of the 2024 International Conference on Advanced Visual Interfaces (AVI ’24). Association for Computing Machinery, New York, NY, USA, 1–3. https://doi.org/10.1145/3656650.3656743
- Ma and Ren (2023) Yumeng Ma and Jiahao Ren. 2023. ProactiveAgent: Personalized Context-Aware Reminder System. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, 1–3. https://doi.org/10.1145/3586182.3625115
- Mackert et al. (2016) Michael Mackert, Amanda Mabry-Flynn, Sara Champlin, Erin E Donovan, and Kathrynn Pounders. 2016. Health Literacy and Health Information Technology Adoption: The Potential for a New Digital Divide. Journal of Medical Internet Research 18, 10 (Oct. 2016), e264. https://doi.org/10.2196/jmir.6349
- Meyers et al. (2020) Nicole Meyers, Alexander F. Glick, Alan L. Mendelsohn, Ruth M. Parker, Lee M. Sanders, Michael S. Wolf, Stacy Bailey, Benard P. Dreyer, Jessica J. Velazquez, and H. Shonna Yin. 2020. Parents’ Use of Technologies for Health Management: A Health Literacy Perspective. Academic Pediatrics 20, 1 (Jan. 2020), 23–30. https://doi.org/10.1016/j.acap.2019.01.008
- Misra et al. (2023) Saumya Misra, Rachana Adtani, Yuvraj Singh, Simran Singh, and Devanshu Thakkar. 2023. Exploring the Factors Affecting Behavioral Intention to Adopt Wearable Devices. Clinical Epidemiology and Global Health 24 (Nov. 2023), 101428. https://doi.org/10.1016/j.cegh.2023.101428
- MoHFW (2005) MoHFW. 2005. School Health Programme.
- Murnane et al. (2018) Elizabeth L. Murnane, Tara G. Walker, Beck Tench, Stephen Voida, and Jaime Snyder. 2018. Personal Informatics in Interpersonal Contexts: Towards the Design of Technology That Supports the Social Ecologies of Long-Term Mental Health Management. Proc. ACM Hum.-Comput. Interact. 2, CSCW (Nov. 2018), 127:1–127:27. https://doi.org/10.1145/3274396
- Napolitano et al. (2023) Enea Vincenzo Napolitano, Simona Fioretto, Elio Masciari, and Arianna Anniciello. 2023. How Pandemic Affected the Adoption of E-Health Systems. In Proceedings of the 27th International Database Engineered Applications Symposium (IDEAS ’23). Association for Computing Machinery, New York, NY, USA, 94–98. https://doi.org/10.1145/3589462.3589496
- Ning et al. (2023) Emma Ning, Andrea T. Cladek, Mindy K. Ross, Sarah Kabir, Amruta Barve, Ellyn Kennelly, Faraz Hussain, Jennifer Duffecy, Scott L. Langenecker, Theresa Nguyen, Theja Tulabandhula, John Zulueta, Olusola A. Ajilore, Alexander P. Demos, and Alex Leow. 2023. Smartphone-Derived Virtual Keyboard Dynamics Coupled with Accelerometer Data as a Window into Understanding Brain Health: Smartphone Keyboard and Accelerometer as Window into Brain Health. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3544548.3580906
- Noblin et al. (2012) Alice M. Noblin, Thomas T. H. Wan, and Myron Fottler. 2012. The Impact of Health Literacy on a Patient’s Decision to Adopt a Personal Health Record. Perspectives in Health Information Management / AHIMA, American Health Information Management Association 9, Fall (Oct. 2012), 1e.
- Nunes et al. (2015) Francisco Nunes, Nervo Verdezoto, Geraldine Fitzpatrick, Morten Kyng, Erik Grönvall, and Cristiano Storni. 2015. Self-Care Technologies in HCI: Trends, Tensions, and Opportunities. ACM Trans. Comput.-Hum. Interact. 22, 6 (Dec. 2015), 33:1–33:45. https://doi.org/10.1145/2803173
- Okolo et al. (2024) Chinasa T. Okolo, Dhruv Agarwal, Nicola Dell, and Aditya Vashistha. 2024. ”If It Is Easy to Understand Then It Will Have Value”: Examining Perceptions of Explainable AI with Community Health Workers in Rural India. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 71:1–71:28. https://doi.org/10.1145/3637348
- Pai and Alathur (2020) Rajesh R. Pai and Sreejith Alathur. 2020. Determinants of Mobile Health Application Awareness and Use in India: An Empirical Analysis. In Proceedings of the 13th International Conference on Theory and Practice of Electronic Governance (ICEGOV ’20). Association for Computing Machinery, New York, NY, USA, 576–584. https://doi.org/10.1145/3428502.3428589
- Pandya et al. (2020) Bhavesh Pandya, Amir Pourabdollah, and Ahmad Lotfi. 2020. Fuzzy Logic Web Services for Real-Time Fall Detection Using Wearable Accelerometer and Gyroscope Sensors. In Proceedings of the 13th ACM International Conference on PErvasive Technologies Related to Assistive Environments (PETRA ’20). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3389189.3397989
- Pasi (2024) Suyash Pasi. 2024. The Impact of Medical Misinformation on Health Rights in India. https://www.humanrightsresearch.org/post/the-impact-of-medical-misinformation-on-health-rights-in-india.
- Patil et al. (2022) Vathsala Patil, Deepak Kumar Singhal, Nithesh Naik, B. M. Zeeshan Hameed, Milap J. Shah, Sufyan Ibrahim, Komal Smriti, Gaurav Chatterjee, Ameya Kale, Anshika Sharma, Rahul Paul, Piotr Chłosta, and Bhaskar K. Somani. 2022. Factors Affecting the Usage of Wearable Device Technology for Healthcare among Indian Adults: A Cross-Sectional Study. Journal of Clinical Medicine 11, 23 (Nov. 2022), 7019. https://doi.org/10.3390/jcm11237019
- Poole (2013) Erika S. Poole. 2013. HCI and Mobile Health Interventions. Translational Behavioral Medicine 3, 4 (Dec. 2013), 402–405. https://doi.org/10.1007/s13142-013-0214-3
- Pulimeno et al. (2020) Manuela Pulimeno, Prisco Piscitelli, Salvatore Colazzo, Annamaria Colao, and Alessandro Miani. 2020. School as Ideal Setting to Promote Health and Wellbeing among Young People. Health Promotion Perspectives 10, 4 (Nov. 2020), 316–324. https://doi.org/10.34172/hpp.2020.50
- Rapp and Cena (2016) Amon Rapp and Federica Cena. 2016. Personal Informatics for Everyday Life: How Users without Prior Self-Tracking Experience Engage with Personal Data. International Journal of Human-Computer Studies 94 (Oct. 2016), 1–17. https://doi.org/10.1016/j.ijhcs.2016.05.006
- Reis et al. (2024) Moritz Reis, Florian Reis, and Wilfried Kunde. 2024. Influence of Believed AI Involvement on the Perception of Digital Medical Advice. Nature Medicine 30, 7 (July 2024), 1–3. https://doi.org/10.1038/s41591-024-03180-7
- Rooksby et al. (2014) John Rooksby, Mattias Rost, Alistair Morrison, and Matthew Chalmers. 2014. Personal Tracking as Lived Informatics. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’14). Association for Computing Machinery, New York, NY, USA, 1163–1172. https://doi.org/10.1145/2556288.2557039
- Saleem and Jan (2021) Sheikh Mohd Saleem and Shah Sumaya Jan. 2021. Modified Kuppuswamy Socioeconomic Scale Updated for the Year 2021. Indian Journal of Forensic and Community Medicine 8, 1 (March 2021), 1–3. https://doi.org/10.18231/j.ijfcm.2021.001
- Samhale (2022) Khalid Samhale. 2022. The Impact of Trust in the Internet of Things for Health on User Engagement. Digital Business 2, 1 (March 2022), 100021. https://doi.org/10.1016/j.digbus.2022.100021
- Scroggs and Love (2022) Barrett Scroggs and Heather A. Love. 2022. Understanding Sexual and Gender Minority Substance Use through Latent Profiles of Ecological Systems. Substance Abuse 43, 1 (2022), 640–648. https://doi.org/10.1080/08897077.2021.1986769
- Seelam et al. (2024) Ananya Seelam, Arnab Paul Choudhury, Connie Liu, Miyuki Goay, Kalika Bali, and Aditya Vashistha. 2024. ”Fact-checks Are for the Top 0.1%”: Examining Reach, Awareness, and Relevance of Fact-Checking in Rural India. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 56:1–56:34. https://doi.org/10.1145/3637333
- Shajari et al. (2023) Shaghayegh Shajari, Kirankumar Kuruvinashetti, Amin Komeili, and Uttandaraman Sundararaj. 2023. The Emergence of AI-Based Wearable Sensors for Digital Health Technology: A Review. Sensors 23, 23 (Jan. 2023), 9498. https://doi.org/10.3390/s23239498
- Shin et al. (2019) Grace Shin, Mohammad Hossein Jarrahi, Yu Fei, Amir Karami, Nicci Gafinowitz, Ahjung Byun, and Xiaopeng Lu. 2019. Wearable Activity Trackers, Accuracy, Adoption, Acceptance and Health Impact: A Systematic Literature Review. Journal of Biomedical Informatics 93 (May 2019), 103153. https://doi.org/10.1016/j.jbi.2019.103153
- Singh et al. (2022a) Nidhi Singh, Richa Misra, Sonali Singh, Nripendra P. Rana, and Sangeeta Khorana. 2022a. Assessing the Factors That Influence the Adoption of Healthcare Wearables by the Older Population Using an Extended PMT Model. Technology in Society 71 (Nov. 2022), 102126. https://doi.org/10.1016/j.techsoc.2022.102126
- Singh et al. (2022b) Nihaal Singh, Rushikesh Shukla, Sourya Acharya, and Samarth Shukla. 2022b. The Past, Present, and Future of Health Economics in India. Journal of Family Medicine and Primary Care 11, 12 (Dec. 2022), 7513–7516. https://doi.org/10.4103/jfmpc.jfmpc_2266_21
- Stuart et al. (2022) Tucker Stuart, Jessica Hanna, and Philipp Gutruf. 2022. Wearable Devices for Continuous Monitoring of Biosignals: Challenges and Opportunities. APL Bioengineering 6, 2 (April 2022), 021502. https://doi.org/10.1063/5.0086935
- Su (2023) Zhaoyuan Su. 2023. Caring for Children’s Health and Wellbeing Through Understanding and Designing for Health Data Literacy. In Proceedings of the 22nd Annual ACM Interaction Design and Children Conference (IDC ’23). Association for Computing Machinery, New York, NY, USA, 785–788. https://doi.org/10.1145/3585088.3593922
- Sulz et al. (2024) Lauren Sulz, Daniel B. Robinson, Hayley Morrison, Josh Read, Ashley Johnson, Lucinda Johnston, and Kim Frail. 2024. A Scoping Review of K–12 Health Education in Canada: Understanding School Stakeholders’ Perceptions. Curriculum Studies in Health and Physical Education 0, 0 (Jan. 2024), 1–23. https://doi.org/10.1080/25742981.2024.2311113
- Sun et al. (2023) Shuyi Sun, Alejandra Ruiz, Sima Pirmoradi, and Katia Vega. 2023. BioSparks: Jewelry as Electrochemical Sweat Biosensors with Modular, Repurposing and Interchangeable Approaches. In Adjunct Proceedings of the 2023 ACM International Joint Conference on Pervasive and Ubiquitous Computing & the 2023 ACM International Symposium on Wearable Computing (UbiComp/ISWC ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, 315–320. https://doi.org/10.1145/3594739.3610787
- Tachtler et al. (2021) Franziska Tachtler, Reem Talhouk, Toni Michel, Petr Slovak, and Geraldine Fitzpatrick. 2021. Unaccompanied Migrant Youth and Mental Health Technologies: A Social-Ecological Approach to Understanding and Designing. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–19. https://doi.org/10.1145/3411764.3445470
- Teixeira Lopes and Henriques (2024) Carla Teixeira Lopes and Mariana Henriques. 2024. Unveiling Health Literacy through Web Search Behavior: A Classification-Based Analysis of User Interactions. In Proceedings of the 2024 Conference on Human Information Interaction and Retrieval (CHIIR ’24). Association for Computing Machinery, New York, NY, USA, 1–11. https://doi.org/10.1145/3627508.3638318
- Teixeira Lopes and Ramos (2020) Carla Teixeira Lopes and Edgar Ramos. 2020. Studying How Health Literacy Influences Attention during Online Information Seeking. In Proceedings of the 2020 Conference on Human Information Interaction and Retrieval (CHIIR ’20). Association for Computing Machinery, New York, NY, USA, 283–291. https://doi.org/10.1145/3343413.3377966
- Thakkar et al. (2022) Divy Thakkar, Azra Ismail, Pratyush Kumar, Alex Hanna, Nithya Sambasivan, and Neha Kumar. 2022. When Is Machine Learning Data Good?: Valuing in Public Health Datafication. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3491102.3501868
- Thapa et al. (2023) Sanjit Thapa, Abubakar Bello, Alana Maurushat, and Farnaz Farid. 2023. Security Risks and User Perception towards Adopting Wearable Internet of Medical Things. International Journal of Environmental Research and Public Health 20, 8 (April 2023), 5519. https://doi.org/10.3390/ijerph20085519
- Tholander and Nylander (2015) Jakob Tholander and Stina Nylander. 2015. Snot, Sweat, Pain, Mud, and Snow: Performance and Experience in the Use of Sports Watches. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (CHI ’15). Association for Computing Machinery, New York, NY, USA, 2913–2922. https://doi.org/10.1145/2702123.2702482
- Thulasi et al. (2024) K Thulasi, Sumathi Balakrishnan, Lu Yue, Rattana Monpreeyadee, Syeda Rahmathunnisa Fatima, Mohammad Izhar Shafiq Bin Rosli, Giok Qi Wei, and MB Malarvili. 2024. IoT-Enabled Wearable Prototype: Detecting Signs of Depression. In Proceedings of the 2023 7th International Conference on Computational Biology and Bioinformatics (ICCBB ’23). Association for Computing Machinery, New York, NY, USA, 78–84. https://doi.org/10.1145/3638569.3638580
- Tong et al. (2019) Catherine Tong, Matthew Craner, Matthieu Vegreville, and Nicholas D. Lane. 2019. Tracking Fatigue and Health State in Multiple Sclerosis Patients Using Connnected Wellness Devices. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 3, 3 (Sept. 2019), 106:1–106:19. https://doi.org/10.1145/3351264
- Ugargol and Bailey (2018) Allen Prabhaker Ugargol and Ajay Bailey. 2018. Family Caregiving for Older Adults: Gendered Roles and Caregiver Burden in Emigrant Households of Kerala, India. Asian Population Studies 14, 2 (May 2018), 194–210. https://doi.org/10.1080/17441730.2017.1412593
- Upadhyay et al. (2023) Nitin Upadhyay, Aakash Kamble, and Ajinkya Navare. 2023. Virtual Healthcare in the New Normal: Indian Healthcare Consumers Adoption of Electronic Government Telemedicine Service. Government Information Quarterly 40, 2 (April 2023), 101800. https://doi.org/10.1016/j.giq.2022.101800
- Volpato et al. (2021) Lucia Volpato, María del Río Carral, Nicolas Senn, and Marie Santiago Delefosse. 2021. General Practitioners’ Perceptions of the Use of Wearable Electronic Health Monitoring Devices: Qualitative Analysis of Risks and Benefits. JMIR mHealth and uHealth 9, 8 (Aug. 2021), e23896. https://doi.org/10.2196/23896
- Wang et al. (2019) Collin Wang, Carolyn Pang, Karyn Moffatt, Rock Leung, and Joanna McGrenere. 2019. Supporting Older Adults in the Use of Smart Devices for Personal Health Management. In Proceedings of the 21st International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’19). Association for Computing Machinery, New York, NY, USA, 615–617. https://doi.org/10.1145/3308561.3354628
- Wani et al. (2024) Asra Sakeen Wani, Ishika Joshi, Nadia Ishfaq Nahvi, and Pushpendra Singh. 2024. ”Unrest and Trauma Stays with You!”: Navigating Mental Health and Professional Service-Seeking in Kashmir. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3613904.3642507
- Wisetsri and Vijai (2021) Worakamol Wisetsri and C. Vijai. 2021. Rise of Artificial Intelligence in Healthcare Startups in India. Advances in Management 14, 1 (March 2021), 48–52.
- Wozniak-O’Connor et al. (2024) Vaughan Wozniak-O’Connor, Megan C. Rose, and Deborah Lupton. 2024. Health Information Ecologies: An Arts-based and Multisensory Approach. In Proceedings of the 35th Australian Computer-Human Interaction Conference (OzCHI ’23). Association for Computing Machinery, New York, NY, USA, 706–711. https://doi.org/10.1145/3638380.3638436
- Zaher et al. (2024) Mahmoud Zaher, Nabil Eldakhly, and Yahia Hassan. 2024. The Emerging Role of Wearable Health Technologies in Proactive Disease Prevention. International Journal of Wireless and Ad Hoc Communication 8 (Jan. 2024), 40–50. https://doi.org/10.54216/IJWAC.080105
- Zegura et al. (2018) Ellen Zegura, Carl DiSalvo, and Amanda Meng. 2018. Care and the Practice of Data Science for Social Good. In Proceedings of the 1st ACM SIGCAS Conference on Computing and Sustainable Societies (COMPASS ’18). Association for Computing Machinery, New York, NY, USA, 1–9. https://doi.org/10.1145/3209811.3209877
- Zhan et al. (2024) Xiao Zhan, Noura Abdi, William Seymour, and Jose Such. 2024. Healthcare Voice AI Assistants: Factors Influencing Trust and Intention to Use. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 62:1–62:37. https://doi.org/10.1145/3637339
- Zhang and Shahriar (2020) Chi Zhang and Hossain Shahriar. 2020. The Adoption, Issues, and Challenges of Wearable Healthcare Technology for the Elderly. In Proceedings of the 21st Annual Conference on Information Technology Education (SIGITE ’20). Association for Computing Machinery, New York, NY, USA, 50–53. https://doi.org/10.1145/3368308.3415454
- Zhang et al. (2022) Zhenming Zhang, Enjun Xia, and Jieping Huang. 2022. Impact of the Moderating Effect of National Culture on Adoption Intention in Wearable Health Care Devices: Meta-analysis. JMIR mHealth and uHealth 10, 6 (June 2022), e30960. https://doi.org/10.2196/30960
- Zhao et al. (2024) Yuchen Zhao, Tulika Banerjee, Na Liu, and Jennifer G Kim. 2024. Grow With Me: Exploring the Integration of Augmented Reality and Health Tracking Technologies to Promote Physical Activity. In Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, 1–9. https://doi.org/10.1145/3613905.3650907